1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

The Banshee (An Approach)

La lupfan- tomo Le loup-garou

The Werewolf

tradukita de Max Knight tradukita de Bertilo Wennergren tradukita de R. Platteau tradukita de Alexander Gross

One night, a banshee

slunk away from mate and child,

and in the gloom went to a village tea-

cher’s tomb, requesting him: “In-

flect me, pray.”

Jen lupfantom’

de famili’ meznokte iam

kuris for al tombo de la

profesor’. kaj petis pri kon-

jugaci’.

Un loup-garou, une cer-

taine nuit, Laissa sa moitié et puis se

rendit Auprès de la tombe d’un

maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

A Werewolf, troubled by

his name, Left wife and brood one night and came To a hidden graveyard to

enlist The aid of a long-dead

philologist.

The village teacher climbed up straight

upon his grave stone

with its plate and to the apparition

said who meekly knelt be-

fore the dead:

Do diktis tiu

dokta ul’ starante sur la

kructabul’ dum lup’ senmo-

va kiel splinto

atentis vortojn de l’ mortinto.

Un loup-garou, une cer-

taine nuit, Laissa sa moitié et puis se

rendit Auprès de la tombe d’un

maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.” The seer arose a bit un-

steady Yawned twice, wheezed once, and then was ready.

“The banSHEE, in the subject’s place;

the banHERS, the

possessive case. The banHER, next,

is what they call objective case–and

that is all.”

De “lupfantomo”, simpla rekta,

tra “lupfintomo”,

plej perfekta, plu “lupfonto-

mo”, iom rara,

ĝis “lupfuntomo”, ho, koŝmara.

“ - On peut à ton sujet puisque nous y sommes Se demander encor bien des choses en somme. Loup-garoù ? mais avant Il faut se dire : loup-

garquand ?

“Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

…  …  …  …

2/3

The banshee marve-

led at the cases and writhed with

pleasure, making faces,

but said: “You did

not add, so far, the plural to the sin-

gular!”

Ĉi participoj plaĉis multe al la fantomo, kaj rezulte ĝi petis: “Legu plu el PIV-o pri participoj en pasivo.”

Loup-garcomment ? et puis Bien évidemment loup-

garqui ? ” Toutes ces questions semblèrent au loup bien absconses,

Car en fait ce qu’il voulait c’était les réponses.

“I know that–I’m no mental cripple– The future form and

participle Are what I crave,” the beast replied. The scholar paused– again he tried:

The teacher, though,

admitted then that this was not wi-

thin his ken. “While bans are fre-

quent”, he advised,

“A she cannot be plurized.”

La viro ĉagreniĝis tamen: “Jam fermis mi per fin’ kaj amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

“ - Point de réponses, dit le puits de savoir, Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes. “ - Oh!, même les baisers de ma douce moitié “A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Hasbeen-wolf?’ Utterly wrong!

Such words are wounds beyond all suture– I’m sorry, but you have

no future.”

The banshee, rising

clammily, wailed: “What about

my family?” Then, being not a

learned creature, said humbly

“Thanks” and left the teacher.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulu-

la.

Ne pourrons plus me faire me sentir tout entier. Rester un loup-garou est un sort qui m’alarme. ” Mais comme le loup était

bien poli Il remercia et puis repartit.

The Werewolf knew better–his sons still slept

At home, and homewards now he crept, Happy, humble, without

apology For such folly of philolo-

gy.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu  ĉi  angli-
go  troviĝas  en

http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en Esperanton de Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu la retejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .